A plane is travelling at 200 mi/h with respect to the air in the direction of the wind. The wind is blowing at 35 mi/hr. A) What is the velocity of the plane with respect to the Earth? B) A passenger on the plane walks from the front to back at 3 mi/hr with respect to the plane. What is the passenger’s velocity with respect to the Earth?
Answer #1

Velocity of plane with respect to earth = velocity of plane with respect to air + velocity of air/wind
V(plane) = V(pa) + Va

Plane is flying in the direction of wind, so sign of both, V(pa) and Va is same.

V(plane) = 200 + 35 = 235 mi/h

Velocity of passenger with respect to earth = velocity of passenger with respect to plane + velocity of plane
V(passenger) = V(pp) + V(plane)
As passenger walks from front to back of plane, its velocity with respect to plane is in the diection opposite to velocity of plane.
V(passenger) = - 3 + 235 = 232 mi/h